The most recent financial statements for Schenkel Co. are shown here: Income Statement Balance Sheet $16,800 25,250 Sales $14,600 Current assets $12,300 Debt Costs Taxable income Taxes (40%) 8,600 Fixed assets 29,750 Equity $42,050 Total $42,050 6,000 2,400 $ 3,600 Total Net income Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 40 percent dividend payout ratio. No external equity financing is possible. What is the sustainable growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Sustainable growth rate Hints References eBook & Resource Hint #1 heck
Answer #1

sustainable growth rate= (ROE × b) / [1 - (ROE × b)]

ROE=Return on Equty

The plowback ratio, b

ROE = NI / TE

=3600/25250=14.26%

b = 1 - .40=.60

Sustainable growth rate = (ROE × b) / [1 - (ROE × b)]
Sustainable growth rate = [.1426*(.60)] / [1 - .1426(.60)]=9.35%